OpenStudy (anonymous):

George makes large cookies. These cookies are 18 inches in diameter. How much icing will he need to cover the top of each cookie? Answer 9π in2 18π in2 81π in2 324π in2

OpenStudy (mertsj):

81 pi

OpenStudy (anonymous):

A=pi*r^2, so A=pi*9^2...so ur answer is 81pi in^2